INAF queue observing, 2015 Nov 06-Nov 07

Observer: S. Bisogni, R. Carini
Telescope Operator: Steve Allanson
Support Astronomer: Dave Thompson on call
Instrument: LBC-MODS

Summary

23:51 The TO is opening the chamber very early so to ventilate and have good temperature for the beginning of the night. At the sunset the sky is clear, the humidity is only 13%

Night duration (12\xB0 el twilight) 19:00-5:10 ( 02:00-12:10 UT) = 10 h 10m

Observing time: 9 hrs 10 m (90%)

Weather time loss: 0 h (0%)

technical time loss 1 h (10%) (problem with MODS2 acquisition)

Data Summary

Project_RUN Objects Filter ExpTime Notes
standard S112 g B, r V    
Clementini PegasusIII, PiscesII B, V 0.5 ob from number 6 to 10
Bellazzini O 339.0+09.0 g,r 0.2 h completed
Bellazzini Q 342.1+20.6 g,r, 0.2 h completed
Bellazzini P 341.7+07.7 g,r 0.2h completed
Saracco M003XLSms1 red grating 2.5h  
Standard gd71 red grating 10min  
Bellazzini A 063.7+33.3 g,r 0.2h completed
Bellazzini E 092.1+09.5 g,r 0.2 h completed
Bellazzini B 084.4+24.0 g,r 0.2 h completed
Bellazzini C 086.4+31.8 g,r 0.2 h completed
Bellazzini D 090.9+30.9 g,r 0.2 h completed
Bellazzini F 100.0+36.7 g,r 0.2 h completed
Bellazzini G 100.9+09.2 g,r 0.2h during the sunset
Skyflat LBC U_spec B, V i

Details

23:51 The TO is opening the chamber very early so to ventilate and have good temperature for the beginning of the night.

00:36 starting running LBC SkyFlat test for S-Y filters, temperature 1.7\xB0, humidity 13%
  • 00:50 running S-Y SlyFlat, counts are low (~10000 in blue and ~ 6000 in red)
  • 00:54 trying test for g-i flats, g perfect counts, i saturated
  • 00:57 running SkyFlats g-i, i counts good now, g a little bit low

01:07 slewing towards photometric standard SA112
  • 01:15 trying running DOFPIA, but the TO says it's still too early
  • 01:22 DIMM seeing 0.6''
  • 01:26 TO tries the co-pointing in the photometric standard field

01:36 run the photometric standard SA112

01:45 slew to Clementini program
  • Pegasus copointing and dofpia
  • dofpia is good, problems with ds9 and iraf
  • do again the copointing

2:01 start science obs with Pegasus_6, then Pisces_6 up to Pegasus_10 and Pisces_10
  • we run dofpia, evey 3 obs.
  • 02:31 seein 0.8''
  • images of Pisces_7 in red a little bit elongated
  • 3:00 seeing 0.8''

3:06 slew to Bellazzini program: target O
  • dofpia and copointing : run 2015B_1_O_pointing.ob
  • 03:17 run the science ob: 2015B_1_O.ob
  • seeing 0.8''-1.0''

3:32 slew to target P
  • dofpia and copointing: run 2015B_1_P_pointing.ob
  • problems with the copointing

*3:45 slew to target Q
  • dofpia and copointg run 2015B_1_Q_pointing.ob
  • 04:00 run the science ob : 2015B_1_Q.ob
  • 4:09 seeing 0.75''

4:15 slew to target P again
  • The field of P is very close to Q, so is not necessary the copointing, we run just the dofpia
  • 4:19 run the science: 2015B_1_P.ob
  • seeing 0.65'''

04:33 switch to MODS

04:57 slew to Saracco program
  • acqMODS2 mods.3.XLSms1.acq mods2r.20151107.0002-0007
  • we should change the guide star because the shadow is very close to the first slit, but this is the only available

WARNING!!!!!!!! the designs of AGw of MODS1 and MODS2 are different, in this particular case the shadow of the patrol field is very close to one slit and we could not change the guide star to resolve the problem because it was the only available. However, there is faint spectrum in this slit (it is the first in the bottom of the image).

  • problem with the MODS2Align: Error: the number of the aligment stars and boxes are not equal
  • problem resolved. we had to remove _modsalign.* generated after Control c during the first MODS2Alingn.

05:52 execMODS2 mods.3XLSms1_4.ob mods2r.20151107.0008-011
  • seeing 0.9''

07:02 execMODS2 mods.3XLSms1_5.ob mods2r.20151107.0012-0015 * seeing 0.9''

08:13 execMODS2 mods.3XLSms1_6.ob (only 2 frames) mods2r.20151107.0016-0017 * seeing 1.1''

08:50 acqMODS2 gd71.acq (Standard Saracco)
  • execMODS2 gd71.obs mods2r.20151107.0020-0022

09:15 back to LBC

09:30 slew to Bellazzini program, target A:
  • dofpia and copointing
  • 09:45 run the science: 2015B_1_A.ob
  • seeing 0.5-0.6''

10:00 slew to the target E
  • dofpia and copointing
  • 10:10 run the science: 2015B_1_E.ob
  • seeing 0.6''

10:25 slew to the target B
  • dofpia and copointing *10.35 run science: 2015B_1_B.ob
  • seeing 0.5''

10:50 slew to the target C
  • dofpia and copointg
  • 10:59 run science : 2015B_1_C.ob
  • seeing 0.5''-0.6''

11:13 slew to the target D
  • dofpia and copointing
  • 11:24 run science 2015B_1_D.ob
  • seeing 0.7''-0.8''

11:39 slew to the target F
  • dofpia and copointing
  • 11.54 run science 2015B_1_F.ob
  • seeing 0.6''-0.7'''

12:10 slew to the target G
  • dofpia and copointing
  • 12:22 run Science : 2015B_1_G.ob
  • seeing 0.5''-0.6''

12:37 flat field , test in VI
  • problems in getting in red images.
  • 12:55 start again the flat fields: some SY are saturated

Article text.

-- %USERSIG{SusannaBisogni - 2015-11-06}%

Comments

 
Topic revision: r12 - 09 Nov 2015, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback